300 apples r distributed equally among a certain no of stdudents had there been 10 more students each would have recd 1 apple less find number of students.
Answer by stanbon(75887)

300 apples r distributed equally among a certain no of stdudents had there been 10 more students each would have recd 1 apple less find number of students.
--------
Let the original # of students be "x";
300/x = k
300/(x+10) = k-1
------
Substitute for "k" and solve for "x":
---
300/(x+10) = (300/x)-1
----
300/(x+10) = (300-x)/x
----
300x = 300x-x^2+3000-10x
-------
-x^2-10x+3000 = 0
---
Positive solution:
x = 50 (original # of students)
==================================
Cheers,
Stan H.
======================